OpenStudy (anonymous):

Solving equation by addition method y=2x-3 4x+4y=-1 how would i solve

OpenStudy (anonymous):

its linear equations

OpenStudy (anonymous):

x=11/12 just use direct subsitution

OpenStudy (anonymous):

4x+4(2x-3) = -1 4x+8x-12 = -1 12x -12 = -1 12x = 11 x= 11/12